A random sample of 150 workers with children in day care shows a mean day care cost of Rs. 3700 and a standard deviation of Rs. 600. Verify the department claim that the mean is equal to Rs. 3500 at the 0.05 level of significance. (Hint Z0.05 = 1.96) 


1
Expert's answer
2020-09-01T18:12:58-0400

Here given:

Accept the Null hypothesis, Department claim if ( -1.96<Z<1.96.

Otherwise Reject the Claim.

Here Given

No. of Workers(n) = 150

Standard deviation( "\sigma" ) = 600

significance level ("\alpha" ) = 0.05

Z(0.05) =( -1.96)<Z<1.96

X = 3500


now , Z-value is

Z =( "X-X" bar)/"\sigma"

=(3500-3700)/600

=(-200)/600

Z = (-1/3)=(-0.33) which is also between (-1.96) to (1.96).

So, the claim is correct. Accept the claim.